OpenStudy (anonymous):

Solve for b: 1/t= 1/a + 1/b A) b= at/t-a B) b= at/a-t Explain.

OpenStudy (anonymous):

1/t= 1/a + 1/b 1/t-1/a=1/b 1/b=(a-t)/at

OpenStudy (anonymous):

did u understand what I did?

OpenStudy (anonymous):

the only thing i am wondering is why is it a-t and not t-a

OpenStudy (anonymous):